THE MOLECULAR STRUCTURES OF 1,2-DISELENIN AND 2-SELENATHIIN J. Z. GILLIES, Department of Chemistry, Siena College, Loudonville, NY 12211; \underline{C. W. GILLIES} and C. NATHAN, Department of Chemistry, Rensselaer Polytechnic Institute, Troy, NY 12180.